of the people at sunrise school 40% are female students 5% are staff and the remaining people are male students. There are 270 more students than female students. How many people are at sunrise school

Answer by checkley79(3341)

.40X+.05X+.05X+270=X
.50X+270=X
.50X-X=-270
-,50X=-270
X=-270/-.50
X=540 TOTAL PEOPLE @ SUNRISE SCHOOL.
PROOF:
.40*540+.05*540+.05*540+270=540
216+27+27+270=540
540=540